OpenStudy (anonymous):

Let f(x) = 3x+2 and g(x) = x^2 Part 1 [2 points] Find f(g(x)) Part 2 [2 points] Find g(f(x)) Part 3 [4 points] Use complete sentences to explain any difference in part 1 and part 2.

OpenStudy (anonymous):

I just dont understand what to do first

OpenStudy (anonymous):

f(g(x)) means put g(x) value for every x in f(x) that means.... 3x^2 +2

OpenStudy (anonymous):

With part 1, in the expression f(g(x)), substitute g(x) with what you know about g(x)

OpenStudy (anonymous):

so would it be f(x)=3(x^2)+2 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so f(x)= 3x^2+ 2?

OpenStudy (anonymous):

g(f(x)) means put the value of f(x) in the x value of g(x) that means (3x+2)^2

OpenStudy (anonymous):

YES thats where i get stuck. How do I simplify (3x+2)^2 ??

OpenStudy (anonymous):

yep.... @kakrazz

OpenStudy (anonymous):

the difference is that both the answers may differ

OpenStudy (anonymous):

how do i simplify (3x+2)^2 ??

OpenStudy (anonymous):

is it 3x^2+4 ??

OpenStudy (anonymous):

apply (a+b)^2 formula dude.... donno that???

OpenStudy (anonymous):

No! Ah omg, my teacher doesnt teach me anything. Okay so how do i do that?

OpenStudy (anonymous):

\[3x ^{2}+2 for (i) and 9x ^{2}+12x+4 for (ii)\]

OpenStudy (anonymous):

so whats the final answer. I thought it was simply- square the 3x and 2...

OpenStudy (anonymous):

\[(a+b)^{2}= a ^{2}+b ^{2}+2ab\]

OpenStudy (anonymous):

put 3x for a and 2 for b

OpenStudy (anonymous):

(3x+2)^2 = 3x^2+2^2+2(3x)(2) ??